P 2 4y - ydy 0

What is the value of this integral ?(according to unit circle)
Answer #1

4 Svay yzdy = Sv4-8-32 dy y=0 Now let 4-2=1 dy=du 4 - u2 du Let u=2Sm V = U-aresm (4) de=&callede rt-48nre du - fecerar = ft